Find an equation of the line containing the given pair of points.
(-3,-1) and (-9,-6)
The equation of the line y = ?
Once again I need some serious help with this.

Answer by stanbon(75887)

Find an equation of the line containing the given pair of points.
(-3,-1) and (-9,-6)
---------------
slope = -6--1/-9--3 = -5/-6 = 5/6
---------------------------------------
Form y = mx + b;
You have y,x,m; solve for b:
-1 = (5/6)*-3 +b
b = 3/2
--------------
EQUATION:
y = (5/6)x + (3/2)
-----------------------
Cheers,
Stan H.